# SymMatika: Structure-Aware Symbolic Discovery

### v1

@@ Abstract
Symbolic regression (SR) seeks to recover closed-form mathematical expressions that describe observed data. While existing methods have advanced the discovery of either explicit mappings (i.e., $y=f(\mathbf{x})$) or discovering implicit relations (i.e., $F(\mathbf{x},y)=0$), few modern and accessible frameworks support both. Moreover, most approaches treat each expression candidate in isolation, without reusing recurring structural patterns that could accelerate search. We introduce SymMatika, a hybrid SR algorithm that combines multi-island genetic programming (GP) with a reusable motif library inspired by biological sequence analysis. SymMatika identifies high-impact substructures in top-performing candidates and reintroduces them to guide future generations. Additionally, it incorporates a feedback-driven evolutionary engine and supports both explicit and implicit relation discovery using implicit-derivative metrics. Across benchmarks, SymMatika achieves state-of-the-art recovery rates, achieving 5.1% higher performance than the previous best results on Nguyen, the first recovery of Nguyen-12, and competitive performance on the Feynman equations. It also recovers implicit physical laws from Eureqa datasets up to 100×100$\times$100 × faster. Our results demonstrate the power of structure-aware evolutionary search for scientific discovery. To support broader research in interpretable modeling and symbolic discovery, we have open-sourced the full SymMatika framework.
@@ 1 Introduction
To improve efficiency, neural-guided SR approaches integrate deep learning to bias the search toward promising expressions. Methods such as DSR Petersen et al. (2019), AI Feynman Mundhenk et al. (2021), and NGGPPS Udrescu and Tegmark (2020) use RNNs or Transformers to guide expression generation or predict symbolic transformations. Despite their improved convergence, these methods often require substantial computational and memory resources, limiting their applicability to large-scale or real-time scientific discovery.
@@
In this work, we introduce SymMatika, a unified symbolic regression framework that discovers both explicit and implicit mathematical relations by combining multi-population feedback-based genetic programming with learned structural pattern reuse. Given data $\mathcal{D}_{E}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ for explicit tasks or $\mathcal{D}_{I}=\{(\mathbf{x}_{i})\}_{i=1}^{N}$ for implicit ones, SymMatika iterates over three key phases: (1) feedback-based GP, which adaptively adjusts operation weights based on evolutionary context; (2) structural analysis, which extracts frequent syntactic substructures from high-performing expressions; and (3) population-specific updates to generation parameters, enabling structurally guided tree construction. The algorithm returns the fittest expression across all populations. In experiments on the Nguyen benchmark and Feynman equations, SymMatika outperforms state-of-the-art methods in accuracy and convergence. It also recovers implicit governing equations from experimental data in the Eureqa dataset with up to 100×100$\times$100 × faster, demonstrating both versatility and performance.
@@ 2 Related Work
GP has long been a cornerstone of symbolic regression, originating with Koza’s foundational work Koza (1990, 1994), which introduced the idea of evolving expression trees via biologically inspired operations such as mutation and crossover. Modern GP-based SR systems such as Operon Burlacu et al. (2020) and PySR Cranmer (2023) inherit this lineage. Operon uses a steady-state GP model with tournament selection, while PySR employs a multi-population island model and simulated annealing to balance exploration and exploitation during evolution. These frameworks rely on fixed operator probabilities, which can hinder adaptive search in complex landscapes.
@@
Despite its versatility, Eureqa is over 16 years old and struggles on modern benchmarks due to fixed operator schedules and heuristics tailored for low-dimensional problems (originally $\leq 4$ variables). SymMatika preserves Eureqa’s implicit-derivative loss but augments it with improved GP techniques including feedback-driven operator tuning, structural motif reuse, and multi-population coordination. These enhancements enable SymMatika to recover implicit equations up to 100×100$\times$100 × faster and achieves superior performance on explicit SR benchmarks such as Feynman and Nguyen.
@@ 3.1 Problem Setup
We represent symbolic expressions as algebraic trees $\tau$, with internal nodes as unary or binary operators (e.g., $+$, $\times$, $\cos$, $\log$) and leaf nodes as constants or variables. A pre-order traversal of $\tau$ yields the symbolic expression $f$, whose quality is assessed using task-specific fitness metrics.
@@
where $x_{p}$ and $x_{q}$ are variables interdependent with $x_{i}$ and $x_{j}$, respectively. We evaluate all possible such pairings and take the worst-case pairing for evaluation to penalize expressions that perform well only under selective dependencies.
@@ 3.2 Feedback-Based Genetic programming
Traditional subtree crossover introduces large structural changes. To enable finer control during late-stage optimization, we introduce single-node crossover, which swaps only individual nodes of matching type (e.g., binary, operator, variable, constant) between two trees. This allows subtle structural adjustments without disrupting well-performing subtrees.

### v2

@@ Abstract
Symbolic regression (SR) seeks to recover closed-form mathematical expressions that describe observed data. While existing methods have advanced the discovery of either explicit mappings (i.e., $y=f(\mathbf{x})$) or discovering implicit relations (i.e., $F(\mathbf{x},y)=0$), few modern and accessible frameworks support both. Moreover, most approaches treat each expression candidate in isolation, without reusing recurring structural patterns that could accelerate search. We introduce SymMatika, a hybrid SR algorithm that combines multi-island genetic programming (GP) with a reusable motif library inspired by biological sequence analysis. SymMatika identifies high-impact substructures in top-performing candidates and reintroduces them to guide future generations. Additionally, it incorporates a feedback-driven evolutionary engine and supports both explicit and implicit relation discovery using implicit-derivative metrics. Across benchmarks, SymMatika achieves state-of-the-art recovery rates on the Nguyen and Feynman benchmark suites, an impressive recovery rate of 61% on Nguyen-12 compared to the next best 2%, and strong placement on the error-complexity Pareto fronts on the Feynman equations and on a subset of 57 SRBench Black-box problems. Our results demonstrate the power of structure-aware evolutionary search for scientific discovery. To support broader research in interpretable modeling and symbolic discovery, we have open-sourced the full SymMatika framework.
@@ 1 Introduction
To improve efficiency, neural-guided SR approaches integrate deep learning to bias the search toward promising expressions. Methods such as DSR Petersen et al. (2019), AI Feynman Mundhenk et al. (2021), and NGGPPS Udrescu and Tegmark (2020) use RNNs or Transformers to guide expression generation or predict symbolic transformations. Deep learning and GP-based approaches are alternative methods of symbolic regression. We make comparisons between GP-based and deep-learning algorithms on benchmark performances later in Sec. 4.
@@
In this work, we introduce SymMatika, a unified symbolic regression framework that discovers both explicit and implicit mathematical relations by combining multi-population feedback-based genetic programming with learned structural pattern reuse. Given data $\mathcal{D}_{E}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ for explicit tasks or $\mathcal{D}_{I}=\{(\mathbf{x}_{i})\}_{i=1}^{N}$ for implicit ones, SymMatika iterates over three key phases: (1) feedback-based GP, which adaptively adjusts operation weights based on evolutionary context; (2) structural analysis, which extracts frequent syntactic substructures from high-performing expressions; and (3) population-specific updates to generation parameters, enabling structurally guided tree construction. The algorithm returns the fittest expression across all populations. In experiments on the Nguyen benchmark and Feynman equations, SymMatika outperforms state-of-the-art methods in accuracy and convergence. It also recovers implicit governing equations from experimental data in the Eureqa dataset with up to $100\times$ faster, demonstrating both versatility and performance.
@@ 2 Related Work
GP has long been a cornerstone of symbolic regression, originating with Koza’s foundational work Koza (1990, 1994), which introduced the idea of evolving expression trees via biologically inspired operations such as mutation and crossover. Modern GP-based SR systems such as Operon Burlacu et al. (2020) and PySR Cranmer (2023) inherit this lineage. Operon uses a steady-state GP model with tournament selection, while PySR employs a multi-population island model and simulated annealing to balance exploration and exploitation during evolution. These frameworks rely on fixed operator probabilities, which can hinder adaptive search in complex landscapes. Feedback-based adaptive crossover-rate in evolutionary computation Guan et al. (2024) proposes a modifiable crossover distribution to optimize crossover points, however it does not control the probability of crossover being selected over other genetic operations (i.e. mutation or single-node crossover). Self-adjusting mutation rates with provably optimal success rules Doerr et al. (2019) proposes updating mutation rates based on fitness-success of offspring, yet it does not consider the overall evolutionary progress of the algorithm.
@@
Despite its versatility, Eureqa is over 16 years old and struggles on modern benchmarks due to fixed operator schedules and heuristics tailored for low-dimensional problems (originally $\leq 4$ variables). SymMatika preserves Eureqa’s implicit-derivative loss but augments it with improved GP techniques including feedback-driven operator tuning, structural motif reuse, and multi-population coordination. These enhancements enable SymMatika to recover implicit equations up to $100\times$ faster and achieves superior performance on explicit SR benchmarks such as Feynman and Nguyen.
@@ 3.1 Problem Setup
We represent symbolic expressions as algebraic trees $\tau$, with internal nodes as unary or binary operators (e.g., ++, $\times$, $\cos$, $\log$) and leaf nodes as constants or variables. A pre-order traversal of $\tau$ yields the symbolic expression $f$, whose quality is assessed using task-specific fitness metrics.
@@
where $x_{p}$ and $x_{q}$ are variables interdependent with $x_{i}$ and $x_{j}$, respectively. $\Delta$ represents finite differences in numerical analysis (including time-series data), or the numerical approximation of the derivative, and $\partial$ represents the calculus partial derivative of a symbolic expression. We evaluate all possible such pairings and take the worst-case pairing for evaluation to penalize expressions that perform well only under selective dependencies.
@@ 3.2 Feedback-Based Genetic programming
Traditional subtree crossover introduces large structural changes. To enable finer control during late-stage optimization, we introduce single-node crossover, which swaps only one same-type node between trees (i.e. binary operation $\leftrightarrow$ binary operation, unary operation $\leftrightarrow$ unary operation, variable $\leftrightarrow$ variable, etc.).
@@
Feedback-based operator scheduling.
Here, $m_{i}$ and $m_{f}$ are the initial and final operator probabilities for each genetic operation (e.g. crossover, mutation), with signs determined by whether each operator frequency should increase or decrease over time. Crossover starts with probability 60% and decreases to 5%, single-node crossover starts with probability 10% and increases to 15%, mutation starts with probability 30% and increases to 80%. This function is safe-guarded by these probability bounds and our probabilistic selection model, so it is resilient to abnormal degradations in average fitness.
Island populations with migration.
3.3 Motif Library and Structure Reuse
After each iteration of GP, we generate new candidates to replace subsets in each population $P$ of worst-performing individuals. Candidate generation in SymMatika is modeled as a parameterized distribution over expression trees, denoted $p(\tau\mid\theta)$, where $\tau$ is a symbolic expression and $\theta$ encodes node-level generation probabilities (e.g., for selecting constants, variables, and operators) Mundhenk et al. (2021); Petersen et al. (2019). These parameters are updated based on frequency feedback from high-fitness individuals: for each operator type $\theta_{i}$, we compute its frequency $f_{\theta_{i}}$ across top-performing expressions, and apply the update:
$\theta_{i}\leftarrow\theta_{i}+\beta_{\theta_{i}}f_{\theta_{i}}$
Here, $\beta_{\theta_{i}}$ is a scaling factor that accounts for natural biases, e.g., binary operators ($\beta_{\theta_{i}}=0.2$) such as ++ and $\times$ may occur more often than unary functions ($\beta_{\theta_{i}}=1.0$) like $\log$ or $\exp$, and are thus adjusted more conservatively. These biases were selected from observations and tuning of our parameterized candidate generator during initial testing. This feedback mechanism enables population-specific adaptation of the generative distribution, which can bias expression construction toward promising token-level patterns.
Initialization and Data Structure.
To implement structural reuse, we maintain a central motif library $\mathbb{M}_{f}$ that stores high-impact symbolic subexpressions across populations. This library is structured as a $d\times k$ table, where $d$ corresponds to the number of data variables and each row holds up to $k$ motifs associated with a particular variable, where $k$ is typically a small integer ($\leq 20$). Each entry in the table consists of a motif $\mathcal{M}_{v}=(\tau\textquoteright,I(\tau\textquoteright))$, where $\tau\textquoteright$ is a symbolic subtree and $I(\tau\textquoteright))$ is its estimated impact score on fitness. We calculate impact score with the following equation $I(\tau\textquoteright))=\mathcal{L}(\tau)-\mathcal{L}(\tau-\tau\textquoteright)$, where $\mathcal{L}(\tau)$ is the loss of the original tree expression and $\mathcal{L}(\tau-\tau\textquoteright)$ is the loss of the original tree without the subtree $\tau\textquoteright$. This quantifies the importance of the substructure to the overall quality of the expression.
Motif Generation.
After each generation, we examine the top $M$ individuals from each island population to identify candidate motifs. For a given expression $\tau$ (e.g., $\tau=3x^{2}-\cos(y)$), we extract the left and right subtrees of each internal node (e.g., $\tau_{L}=3x^{2}$, $\tau_{R}=-\cos(y)$. Note: $\cos(y)$ is a negative term in the example, so we extract it as $\cos(y)$) and assess their fitness contribution.
Structure-Aware Expression Synthesis.
Once the motif library is updated, new expressions are synthesized by sampling one or more motifs from each row of $\mathbb{M}_{f}$ until all variables are represented. These motifs are then assembled into full expressions using randomly selected binary algebraic operators such as ++, -, $\times$, or $\div$, resulting in composite symbolic expressions $\tau\textquoteright$. The resulting expressions are inserted into a dedicated motif population $P_{\mathcal{M}}$, which co-evolves alongside the main island populations.
Co-Evolve Motif-Population and Main Populations.
3.4 Implementation Details
4 Experiments
For all experiments, SymMatika is instantiated with population sizes of 10,000 individuals per island, followed by truncation to the top 400 for the GP-loop. The number of islands $I$ is set proportional to expression complexity as $I\propto 2d$, where $d$ is the number of variables (capped at $I=8$ for large SR-problems with five or more variables due to computational considerations). We run each experiment for 1500 generations of GP.
We report: (1) recovery rates on the Nguyen benchmark suite (with an emphasis on Nguyen-12 recoveries), (2) error-complexity plots of the Feynman equations (mean proportion of $R^{2}>0.99$ vs. equation complexity) and 57 SRBench Black-box problems (median $R^{2}$ vs. equation complexity), and (3) implicit-relation discovery in physical systems using Eureqa datasets. We also performed ablation studies to assess the contribution of core components of SymMatika. Our goals are to demonstrate the effectiveness of SymMatika on both explicit and implicit SR tasks, highlight cases where it outperforms prior work, and understand how different algorithm modules contribute to performance.
All experiments are run on a 2023 Apple M3 Max MacBook Pro with a 14-core CPU, 30-core GPU, and 36GB unified memory. Unlike other neural-guided approaches, SymMatika does not require GPUs nor access to large language models (LLMs). Since SymMatika uses multi-core execution, and many baselines are either single-core or multi-core with unknown runtimes, we report recovery rates (not wall-clock time) for Nguyen and Feynman.
4.1 Nguyen Benchmark
Table 1: Exact recovery rates (%) on the Nguyen benchmark over 100 runs. SymMatika achieves the highest average performance (96.5%) and recovers Nguyen-12 at significantly higher rates than existing algorithms (61% success).
Dataset Expression SymMatika NGGPPS DSR PQT Eureqa Operon PySR
Nguyen-1 $x^{3}+x^{2}+x$ 100 100 100 100 100
Nguyen-2 $x^{4}+x^{3}+x^{2}+x$ 100 100 100 99 100
Nguyen-3 $x^{5}+x^{4}+x^{3}+x^{2}+x$ 100 100 100 86 95
Nguyen-4 $x^{6}+x^{5}+x^{4}+x^{3}+x^{2}+x$ 100 100 100 93 70
Nguyen-5 $\sin(x^{2})\cos(x)-1$ 100 100 72 73 73
Nguyen-6 $\sin(x)+\sin(x+x^{2})$ 100 100 100 98 100
Nguyen-7 $\log(x+1)+\log(x^{2}+1)$ 98 97 35 41 85
Nguyen-8 $\sqrt{x}$ 100 100 96 21 0
Nguyen-9 $\sin(x)+\sin(y^{2})$ 100 100 100 100 100
Nguyen-10 $2\,\sin(x)\cos(y)$ 100 100 100 91 64
Nguyen-11 $x^{y}$ 100 100 100 100 100
Nguyen-12 $x^{4}-x^{3}+\tfrac{1}{2}y^{2}-y$ 61 0 0 0 0 2 0
Average 96.5 91.4 83.6 75.2 73.9
Figure 3: Error–complexity (rank) trade-off on the Black-box (left) and Feynman (right) datasets.
Tab. 1 reports recovery rates over 100 independent runs per task. We take reported results from Petersen et al. (2019); Mundhenk et al. (2021) on NGGPPS (Note: PQT is one method of training NGGPPS’ RNN), DSR, and Eureqa. SymMatika records an average recovery rate of 96.5%, outperforming all other methods by statistically significant margins (maximum $p<10^{-5}$ from $z$-score testing). We ran 100 runs with random seeds on Nguyen-12 using PySR and Operon with the exact experiment settings (details in Appendix A.3) and reported 0/100 and 2/100 recoveries respectively; given the simpler nature of Nguyen 1-11, we felt it was not necessary to perform testing on problems proven to be easily solvable. Notably NGGPPS Mundhenk et al. (2021) reports a 12% success rate on a relaxed variant with expanded data range (Nguyen-$12^{\star}$), our performance remains significantly stronger on the original Nguyen-12 dataset. These results highlight the advantage of combining structural motif reuse with feedback-driven parameter tuning, especially for long and composite expressions.
4.2 Feynman Equations Benchmark
Table 2: Exact recovery counts on the Feynman benchmark (100 tasks). SymMatika recovers 73 expressions, outperforming all other algorithms.
GPlearn AFP-FE DSR uDSR AIFeynman PySR LaSR SymMatika
20/100 26/100 23/100 40/100 38/100 59/100 72/100 73/100
Next, we evaluate on the Feynman benchmark, a widely adopted dataset of 100 symbolic physics expressions derived from the Feynman Lectures. As in prior work, expressions are sampled to form datasets $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, and success is defined as recovery of the symbolic target. We performed two sets of experiments: (1) recovery rates on 100 Feynman problems outlined in LaSR Grayeli et al. (2024), and (2) error-complexity plot (mean $R^{2}>0.99$ proportion vs. equation complexity) against SRBench La Cava et al. (2021) results. We use the same experimental settings as for the Nguyen benchmark (Appendix A.3).
As shown in Tab. 2, SymMatika recovers 73/100 expressions outperforming all methods, including GPlearn Stephens et al. (2016), DSR Petersen et al. (2019), uDSR Landajuela et al. (2022), AIFeynman Udrescu and Tegmark (2020), PySR Cranmer (2023), and LaSR Grayeli et al. (2024). LaSR Grayeli et al. (2024) extends PySR Cranmer (2023) with natural language priors from LLMs via a ConceptAbstraction function, enabling semantic guidance during search. While this approach is orthogonal to ours, our results show that structurally guided recombination and operator adaptation alone are competitive with language-informed strategies.
To showcase some sample output, we run SymMatika on a sample equation from the Feynman Lectures on Physics Feynman et al. (2015), the relativistic Doppler shift formula $\omega=\frac{1+\frac{v}{c}}{\sqrt{1-\frac{v^{2}}{c^{2}}}}\omega_{0}$. After running SymMatika on its dataset, it returns the expression $\omega=\sqrt{\frac{c+v}{\left(c-v\right)}}\omega_{0}$ with an MLE loss of $3.10862\times 10^{-16}$, and the discovered expression is mathematically equivalent to the ground truth expression, albeit syntactically slightly different. We provide outputs on a subset of sampled Feynman Equations in Appendix Sec. A.1.
For the error-complexity plot, SymMatika places impressively on the Pareto front. We report mean $R^{2}>0.99$ proportion of 0.942 and complexity of 13.294 across 6 trials (Fig. 3). We surpass all algorithms on SRBench – including Operon, AIFeynman and SBG-GP to name a few — with maximum $p=0.044$ — except for MRGP Arnaldo et al. (2014), which reports mean proportion 0.931 yet complexity 26316, significantly larger than SymMatika. On the Black-box subset, however, we report much better results in both error and complexity then MRGP. Additionally, two versions of TPSR Shojaee et al. (2023) slightly outperform our mean ($R^{2}>0.99$) accuracy with 0.949 (TPSR[$\lambda=0.1$]) and 0.952 (TPSR[$\lambda=0$]) to our 0.942, however our complexity is significantly improved with our 13.294 compared to their 84.42 (TPSR[$\lambda=0$]) and 57.22 (TPSR[$\lambda=0.1$]). Overall, SymMatika places strongly on the Pareto front and leads in model complexity.
4.3 Black-box Benchmark
We evaluated SymMatika on 57 Black-box problems gathered from SRBench used by TPSR Shojaee et al. (2023) which include synthetic and real-world datasets and we recorded median $R^{2}$ and mean equation complexity for the error-complexity plot (Fig. 3). We ran 5 full trials and recorded median $R^{2}$ of 0.931 and equation complexity of 28.059. We outperform nearly all benchmarked algorithms in median $R^{2}$ and match TPSR $\lambda=0.5$] (complexity 82.58), where $\lambda$ is a hyperparameter balancing fitting and complexity reward), and falling slightly behind TPSR $\lambda=0$] at 0.938 (complexity 129.85), Operon at 0.937 (complexity 61.74) and TPSR $\lambda=0.1$] at 0.945 (complexity 95.71). Our complexity significantly improves on these algorithms by a factor of 3-4.5 for TPSR and a factor of 2 for Operon, and MRGP which reports a much-lower median $R^{2}$ of 0.502 larger and complexity 9878.172. TPSR records one trial so we cannot report meaningful $p$-testing given limited experimental data, yet we can report that the Operon’s outperformance of us is not statistically significant ($p=0.71138$).
4.4 Eureqa Physical Systems
To evaluate implicit symbolic regression, we compared against Eureqa on four time-series physical systems from its supplementary materials, including circular, pendulum, spherical, and double pendulum motions. Unlike the explicit SR benchmarks, these systems do not provide labeled outputs $y$, and recovery involves discovering hidden algebraic constraints $f(\mathbf{x})=0$ from the raw dynamics.
Table 3: Runtime (s) on Eureqa dataset. SymMatika converges $\times100×1$$0\times$ faster.
System SymMatika (s) Eureqa (s)
circle_1 0.15 25
pendulum_h_1 0.98 31
sphere_1 3.00 320
dbl_pendulum_h_1 900.00 27000
[…]
4.5 Ablation Study
Recovery rate (%)
Problem Baseline $\theta$ Generation Motif Library Full Model
Nguyen-1 100 100 100 100
Nguyen-2 100 100 100 100
Nguyen-3 100 100 100 100
Nguyen-4 90 95 100 100
Nguyen-5 95 100 100 100
Nguyen-6 85 90 95 100
Nguyen-7 60 80 85 95
Nguyen-8 100 100 100 100
Nguyen-9 100 100 100 100
Nguyen-10 100 100 100 100
Nguyen-11 100 100 100 100
Nguyen-12 5 20 40 65
Nguyen average 86.3 90.4 93.3 96.7
95% confidence intervals $\pm$15.92 $\pm$13.02 $\pm$9.82 $\pm$5.7
5 Conclusions
6 Limitations
The primary limitation in our approach is our results on the Feynman Equations. Although we report results outperforming or performing on par with leading models, we are still unable to solve a subset of 27 problems. These problems generally involve long expressions with complicated structures, such as $n=\frac{n_{0}}{\exp(\mu_{m}B/(k_{b}T))+\exp(-\mu_{m}B/(k_{b}T))}$, problem $\text{II}.35.18$ in the Feynman Equations Udrescu and Tegmark (2020). Other unsolved problems include large arguments in unary operators (e.g. $x=\sqrt{x_{1}^{2}+x_{2}^{2}-2x_{1}x_{2}\cos(\theta_{1}-\theta_{2})}$). To attempt further discovery of Feynman equations, we will conduct experiments in an improved hardware setup with more CPU cores to evolve more island populations in parallel with larger populations and for longer. Additionally, we will further investigate methods of discovering implicit relations in higher-dimensional data (i.e. $\geq 5$ variables).
7 Acknowledgments
Appendix
A.1 Sample Outputs From Subset of Feynman Equations
Table 5: Comparison of ground-truth vs. discovered equations.
Equation Number Ground Truth Equation Discovered Equation
I.18.4 $\frac{m_{1}r_{1}+m_{2}r_{2}}{m_{1}+m_{2}} r=r$$\frac{r_{1}m_{1}}{\left(m_{1}+m_{2}\right)}+\frac{r_{2}m_{2}}{\left(m_{1}+m_{2}\right)}$
I.37.4 $_{*}=I_{1}+I_{2}+2\sqrt{I_{1}I_{2}}\cos{\left(\delta\right)} I=I$$_{*}=I_{1}+I_{2}+2\sqrt{I_{1}I_{2}}\cos{\left(\delta\right)}$
I.27.6 $_{f}=\frac{d_{2}}{\left(n+\frac{d_{2}}{d_{1}}\right)} ff$$_{f}=\frac{1}{\frac{1}{d_{1}}+\frac{n}{d_{2}}}$
I.47.23 $c=\sqrt{\frac{\gamma pr}{\rho}}$ $c=\sqrt{\frac{\gamma pr}{\rho}}$
I.12.11 $F=q(E_{f}+Bv\sin(\theta))$ $F=qE_{f}+qBv\sin(\theta)$
A.2 SymMatika Algorithm